### v1

# Euler characteristic of crepant resolutions of specific modular quotient singularities

Abstract
In this paper, we consider a generalization of McKay correspondence in positive characteristic regarding Euler characteristic of crepant resolutions given by finite subgroups of the special linear group. As the main result, we prove that the generalization holds for groups with a specific semidirect product structure. Apart from such groups, two certain examples for groups with more difficult structure are also given.
1 Introduction
Euler characteristic of crepant resolutions is one of important geometric invariants in study of McKay correspondence. As a series of results connecting algebra with geometry, McKay correspondence studies relations between algebraic properties of groups and geometric properties of the associated quotient singularities. Over complex numbers, it was conjectured by Reid and proved by Batyrev via motivic integration ([batyrev1999non]), that for a finite group $G\subseteq\mathrm{SL}(n,\mathbb{C})$ and the associated quotient singularity $X:=\mathbb{C}^{n}/G$, assuming existence of a crepant resolution $Y\to X$, the Euler characteristic $e(Y)$ equals the number of conjugacy classes #$\#\mathrm{Conj}(G)$.
In this paper, the case when the base field is of positive characteristic is focused. Let $k$ be an algebraically closed field of prime characteristic $p>0$, and $G\subseteq\mathrm{SL}(n,k)$ be a finite group giving the quotient singularity $X:=\mathbb{A}^{n}_{k}/G$. For a smooth $k$-variety, its (topological) Euler characteristic is defined as the alternating sum of Betti numbers given by $l$-adic cohomology with compact support, where $l$ is a prime different from $p$.
It is natural to ask if the equality between Euler characteristic of crepant resolutions and number of conjugacy classes still holds. If the order of $G$ is not divided by $p$, then there is an affirmative answer known as tame McKay correspondence (the reader can refer to [wood2015mass], Section 5.2). Unfortunately, in modular case, where #$\#G$ is divided by $p$, the analog of Batyrev’s theorem has been disproved by construction of counterexamples, such as Chen, Du and Gao’s counterexample in non-small case ([chen2020modular]), Yamamoto’s construction for the quotient variety given by the symmetric group $S_{3}\subseteq\mathrm{SL}(3)$ in characteristic $3$ ([yamamoto2021crepant]) and the author’s construction for the quotient variety given by permutation action of the alternating group $A_{4}$ in characteristic $2$ ([fan2023crepant]).
On the other hand, there are also several known results implying that the analog of Batyrev’s theorem in positive characteristic holds for specific modular quotient singularities. We list them as below.
Theorem 1.1 (Known results).
  1. 1.
    ([yasuda2014cyclic], Corollary 6.21) Let $k$ be an algebraically closed field of characteristic $p>0$. Suppose $G\subseteq\mathrm{SL}(n,k)$ is a finite group with no reflections, such that $G\cong C_{p}$, where $C_{p}$ is the $p$-cyclic group. If the associated quotient singularity has a crepant resolution $Y\to X:=\mathbb{A}^{n}_{k}/G$, then $e(Y)=p$.
  2. 2.
    ([yamamoto2021crepant], Theorem 1.2) Let $k$ be an algebraically closed field of characteristic $3$. Suppose $G\subseteq\mathrm{SL}(3,k)$ is a finite group with no reflections, such that $G\cong H\rtimes C_{3}$, where $H$ is a non-modular abelian group and $C_{3}$ is the cyclic group of order $3$. Then the associated quotient singularity has a crepant resolution $Y\to X:=\mathbb{A}^{3}_{k}/G$, and $e(Y)=\#\mathrm{Conj}(G)$.
  3. 3.
    ([fanmaster], Corollary 3.1) Let $k$ be an algebraically closed field of characteristic $p=2$. For any positive integer $l$ with $(p,l)=1$, let $\zeta_{l}$ be a primitive $l$-th root of unity in $k$. For integers $a_{1},a_{2},a_{3},a_{4}$, denote the diagonal matrix $\mathrm{diag}(\zeta_{l}^{a_{1}},\zeta_{l}^{a_{2}},\zeta_{l}^{a_{3}},\zeta_{l}^{a_{4}})$ by $\frac{1}{l}(a_{1},a_{2},a_{3},a_{4})$. Suppose $G\subseteq\mathrm{SL}(4,k)$ has a semidirect product structure $G\cong H\rtimes C_{2}$, where $C_{2}$ is generated by the permutation $(12)(34)$ as an element in $\mathrm{SL}(4)$, and $H$ is one of the following:
    • $H=\langle\frac{1}{l}(1,-1,0,0),\frac{1}{l}(1,0,-1,0),\frac{1}{l}(1,0,0,-1)\rangle$,
    • $H=\langle\frac{1}{l_{1}}(1,-1,0,0),\frac{1}{l_{2}}(0,0,1,-1)\rangle$.
    Then $G$ has no reflections, and the associated quotient singularity has a crepant resolution $Y\to X:=\mathbb{A}^{4}_{k}/G$, satisfying $e(Y)=\#\mathrm{Conj}(G)$.
In the results above, $G$ is always small (that is, $G$ has no reflections) and has a semidirect product structure given by a non-modular abelian normal subgroup and a $p$-cyclic group. If $G\cong H\rtimes K$, it is a classical idea to consider the decomposition $\mathbb{A}^{n}\xrightarrow{/H}\mathbb{A}^{n}/H\xrightarrow{/K}\mathbb{A}^{n}/G$ and the singularities given by $H$ and $K$ respectively. Under the assumption, $H$ is non-modular, and the singularity given by $K\cong C_{p}$ is studied by Theorem 1.1.1, with the same statement as Batyrev’s theorem. In fact, Theorem 1.1.2 and Theorem 1.1.3 are parallel to Ito’s construction of crepant resolutions of trihedral singularities ([ito1994crepant]) and specific $4$-dimensional quotient singularities ([hayashi2017existence]) over $\mathbb{C}$, with the same approach. In this paper, we show that for such small groups with semidirect product structures, the analogous statement of Batyrev’s theorem always holds.
Theorem 1.2 (Main theorem).
Let $k=\overline{\mathbb{F}_{p}}$ be the algebraic closure of the finite fields of characteristic $p>0$, and $G\subseteq\mathrm{SL}(n,k)$ be a finite group with no reflections, such that $G$ has a structure of semidirect product as $H\rtimes C_{p}$, where $H$ is a non-modular abelian group and $C_{p}$ denotes the $p$-cyclic group. If $X:=\mathbb{A}^{n}_{k}/G$ has a crepant resolution $f:Y\to X$, then
[…]
Hereby, Theorem 1.1.2 and Theorem 1.1.3 become examples of the main theorem. By this result, we also point out that the number of isomorphic classes of indecomposable $kG$-modules, which is not necessarily equal to #$\#\mathrm{Conj}(G)$ in positive characteristic, is a candidate for the algebraic invariant of $G$ corresponding to Euler characteristic of the crepant resolution. Indeed, we have the following conjecture.
Conjecture 1.3.
Let $k$ be an algebraically closed field of characteristic $p>0$, and $G\subseteq\mathrm{SL}(n,k)$ be a finite small group of finite representation type (that is, there are finitely many isomorphic classes of indecomposable $kG$-modules). If $X:=\mathbb{A}^{n}_{k}/G$ has a crepant resolution $f:Y\to X$, then
[…]
This conjecture is true for all known cases, including non-modular case, the main theorem and Yamamoto’s counterexample in [yamamoto2021crepant]. On the other hand, there are also examples where the group is of infinite representation type, such as the author’s construction in [fan2023crepant]. In the last section, two quotient singularities given by certain groups of infinite representation type are considered.
Example 1.4 (=Corollary 5.5).
Let $k$ be an algebraically closed field of characteristic $2$, $A_{4}$ and $C_{2}^{2}$ be subgroups of $\mathrm{SL}(4,k)$ given by permutation actions of the alternating group and its normal subgroup containing elements of order $2$ respectively. Denote the associated quotient singularities by $X_{1}:=\mathbb{A}^{4}/{C_{2}^{2}}$ and $X_{2}:=\mathbb{A}^{4}/A_{4}$. If there exists a crepant resolution $Y_{1}\to X_{1}$ ($resp.$ $Y_{2}\to X_{2}$), then $e(Y_{1})=6$ ($resp.$ $e(Y_{2})=10$).
This paper is organized as follows. In Section 2, we introduce results and notations used for the proof of the main theorem, including the main tool - mass formula version of wild McKay correspondence. In Section 3, we prove the main theorem when $G$ is abelian. In Section 4, we prove the main theorem when $G$ is non-abelian. In Section 5, we provide two examples of computation of Euler characteristic of crepant resolutions for groups of infinite representation type, using key idea of proof of the main theorem.
Acknowledgements.
The author expresses his gratitude to Prof. Takehiko Yasuda and Dr. Rin Gotou for valuable discussions. The author also thanks his supervisor, Prof. Yukari Ito, for her helpful advice and consideration.
2 Preliminaries
For the proof of our main theorem, wild McKay correspondence as mass formulas is an important tool. In this version of motivic wild McKay correspondence over finite fields, stringy motives are studied by its realization as stringy point-count #$\#_{\mathrm{st}}$. Under specific conditions, if crepant resolutions exist, then some of their geometric properties can be studied via stringy point-count of the quotient varieties. For more details, the reader can refer to [yasuda2017wild]. Here below we only list results that are necessary for our proof.
Definition 2.1.
Let $K$ be a field and $G$ be a finite group. $M$ is called a $G$-étale $K$-algebra, if $M$ is a finite étale $K$-algebra of degree #$\#G$ equipped with a $G$-action, such that $M^{G}=K$. Isomorphisms of $G$-étale $K$-algebras are isomorphisms of $K$-algebras that are $G$-equivariant. $G\text{-}\mathrm{\acute{E}t}(K)$ is the set consisting of isomorphic classes of $G$-étale $K$-algebras.
Remark 2.2.
Different actions of $G$ on the same $K$-algebra may result in different $G$-étale $K$-algebras. However, conjugate endomorphisms of $G$ do not change the class in $G\text{-}\mathrm{\acute{E}t}(K)$. For each $M\in G\text{-}\mathrm{\acute{E}t}(K)$, there exists a subgroup $G_{M}\subseteq G$ (unique up to conjugation) such that $M\cong L_{M}^{\oplus(G:G_{M})}$, where $L_{M}$ is a Galois extension of $K$ with Galois group isomorphic to $G_{M}$.
Theorem 2.3 ([yasuda2017wild], Proposition 8.5).
Let $k$ be a finite field, $\mathbb{F}_{q}/k$ be a field extension with order $q=p^{e}$, and $K=\mathbb{F}_{q}((t))$ be the local field of formal Laurent series. Let $G$ be a finite subgroup of $\mathrm{SL}(n,k)$ with no reflections. Let $X:=\mathbb{A}^{n}/G$ be the associated quotient variety. Then
#$\#_{\mathrm{st},\mathbb{F}_{q}}X=\sum_{M\in G\text{-}\mathrm{\acute{E}t}(K)}\frac{q^{n-\mathbf{v}_{V}(M)}}{\#C_{G}(G_{M})};$
if furthermore $X$ has a crepant resolution $Y\to X$, then
#$\#Y(\mathbb{F}_{q})=\#_{\mathrm{st},\mathbb{F}_{q}}X=\sum_{M\in G\text{-}\mathrm{\acute{E}t}(K)}\frac{q^{n-\mathbf{v}_{V}(M)}}{\#C_{G}(G_{M})}.$
Here #$\#Y(\mathbb{F}_{q})$ counts number of $\mathbb{F}_{q}$-rational points of $Y$, $C_{G}(G_{M})$ is the centralizer of $G_{M}$ in $G$, and $\mathbf{v}_{V}(M)$ is the $\mathbf{v}$ function defined as follows.
Definition 2.4 ([yasuda2017wild]; $\mathbf{v}$ function).
$\mathbf{v}$ function is defined for an arbitrary finite group $G$ with a representation $\rho:G\to\mathrm{GL}(n,k)\hookrightarrow\mathrm{GL}(n,\mathcal{O}_{K})$. Let $V=\mathbb{A}^{n}_{\mathcal{O}_{K}}$ be the representation space and $\mathcal{O}_{K}[x_{1},\dots,x_{n}]$ be its coordinate ring with linear part $T=\sum\mathcal{O}_{K}x_{i}$. Then for any $M\in G\text{-}\mathrm{\acute{E}t}(K)$, $G$ has both an action on $\mathcal{O}_{M}$ via definition of $G$-étale $K$-algebras and an action on $T$ via the representation $\rho$. Denote by $\mathrm{Hom}_{\mathcal{O}_{K}}^{G}(T,\mathcal{O}_{M})$ the homomorphisms that commute with the actions of $G$. Then $\mathrm{Hom}_{\mathcal{O}_{K}}^{G}(T,\mathcal{O}_{M})$ is a free $\mathcal{O}_{K}$-module of rank $n$ ([yasuda2017toward], Proposition 6.3), and
$\mathbf{v}_{V}(M):=\frac{1}{\#G}\mathrm{length}_{\mathcal{O}_{K}}\frac{\mathrm{Hom}_{\mathcal{O}_{K}}(T,\mathcal{O}_{M})}{\mathcal{O}_{M}\mathrm{Hom}_{\mathcal{O}_{K}}^{G}(T,\mathcal{O}_{M})}.$
Remark 2.5.
$\mathbf{v}$ function is convertible and additive ([wood2015mass], Lemma 3.4). We would use following facts: Even if $\rho$ is not faithful, it can be decomposed as $G\rightarrow G/H\rightarrow\mathrm{GL}(n,k)$ for some $H\subseteq G$, and $\mathbf{v}_{V}(M)=\mathbf{v}_{V_{G/H}}(M^{H})$; if $V$ can be decomposed as $V=\bigoplus_{i=1}^{s}V_{i}$, then $\mathbf{v}_{V}=\sum_{i=1}^{s}\mathbf{v}_{V_{i}}$.
Remark 2.6.
The value of $\mathbf{v}$ function for $M\in G\text{-}\mathrm{\acute{E}t}(K)$ is determined by $L_{M}/K$ as a $G_{M}$-étale $K$-algebra. We simply write $\mathbf{v}(L)$ if $G_{M}$ is fixed.
Next, we give several notations and use them to rewrite the wild McKay correspondence such that Euler characteristic of crepant resolutions can be studied.
Definition 2.7.
Denote
[…]
We define two series of $q$, denoted by $f_{G}$ and $F_{G}$, as follows.
$f_{G}(q):=\sum_{L\in A(G)}q^{n-\mathbf{v}(L)},$
$F_{G}(q):=\sum_{[G^{\prime}]\subseteq G}\frac{1}{\#N_{G}(G^{\prime})}f_{G^{\prime}}.$
Here $N_{G}(G^{\prime})$ is the normalizer of $G^{\prime}$ in $G$, and $[G^{\prime}]\subseteq G$ is taken up to conjugation.
Remark 2.8.
Every element in $A(G)$ contains information of a Galois extension $L/K$ with Galois group $\mathrm{Gal}(L/K)\cong G$ and the action of $G$ on it. For example, if $G\cong C_{l}$ for some prime $l$ ($l\neq p$), then by Kummer theory, there are $l+1$ Galois extensions of $K=k((t))$ with $l$-cyclic Galois group up to isomorphism: one is unramified and the other $l$ extensions are totally ramified. On the other hand, for each extension, there are $l-1$ ways to map the given generator of $G$ to a nontrivial element in the Galois group, giving different actions up to isomorphism of $G$-étale algebras. Hence $A(C_{l})$ has $l^{2}-1$ elements.
Remark 2.9.
[…]
Definition 2.10.
Let $f$ be a Laurent polynomial with coefficients in $\mathbb{Q}$. We define $S(f):=f(1)$ as the sum of coefficients of $f$. $S$ is therefore a homomorphism between $\mathbb{Q}$-algebras.
Corollary 2.11.
Under the assumption of Theorem 2.3, if crepant resolution $Y$ exists, then $F_{G}(T^{2})$ converges as a $\mathbb{Z}$-coefficient polynomial in $T$, and the (topological) Euler characteristic $e(Y)=S(F_{G})$.
Proof.
The first assertion is from motivic wild McKay correspondence (construction of stringy motives and their realization as Poincaré polynomials of crepant resolutions via $\mathbb{L}\mapsto T^{2}$), and the second one follows; one can also check it by the functional equation in Weil conjectures. The reader can refer to [yasuda2014cyclic] for details. ∎
Proposition 2.12.
If $G$ is non-modular, then $F_{G}$ is a polynomial in $q$ and $S(F_{G})=\#\mathrm{Conj}(G)$. If $G$ is also abelian, then $\sum_{G^{\prime}\subseteq G}f_{G^{\prime}}=(\#G)^{2}$.
Proof.
This is a corollary of tame McKay correspondence ([wood2015mass], Section 5.2). We also point out that $S(f_{G})=\#A(G)$ here. ∎
Proposition 2.13.
If $G$ is isomorphic to a cyclic group of order $p$ and there exists a crepant resolution $Y\to X$ , then $F_{G}$ is a polynomial in $q$, $S(F_{G})=p$ and $S(f_{G})=p^{2}-1$.
Proof.
This is a corollary of $p$-cyclic McKay correspondence ([yasuda2014cyclic], Corollary 6.21). ∎
In the last part of this section, we give the proof for the equality between number of conjugacy classes and number of isomorphic classes of indecomposable representations in the statement of main theorem.
Proposition 2.14.
Let $G$ be a finite group equipped with a semidirect product structure $H\rtimes C_{p}$, where $H$ is abelian and $C_{p}$ is a cyclic group of prime order $p$. Let $C(G)$ be the center of $G$ and $k$ be a splitting field of $G$ of characteristic $p$ (equivalently, for any element $h\in H$, $k$ contains all of the $\mathrm{ord}(h)$-th primitive roots of unity in $\overline{k}$). Then
#$\#\mathrm{Conj}(G)=\left\{\begin{array}[]{ll}\displaystyle p\#C(G)+\frac{\#H-\#C(G)}{p}&G\text{ is not abelian,}\\ \#G&G\text{ is abelian.}\end{array}\right.$
which coincide the number of isomorphic classes of indecomposable $kG$-modules.
Proof.
We first consider the case when $G$ is not abelian. Then $G$ contains #$\#C(G)$ conjugacy classes of $1$ element, #$\frac{\#H-\#C(G)}{p}$ conjugacy classes of $p$ elements and $(p-1)\#C(G)$ conjugacy classes of $(H:C(G))$ elements. Hence the number of conjugacy classes of $G$ is equal to right hand side in the proposition.
On the other hand, by a corollary of Green correspondence (one can refer to [webb2016course], Corollary 11.6.5), for $G$ with its $p$-Sylow subgroup isomorphic to $C_{p}$, if one denotes number of simple $kG$-modules up to isomorphism by $l_{k}(G)$, then
#$\#\{[\text{indecomposable }kG\text{-modules}]\}=l_{k}(G)+(p-1)l_{k}(N_{G}(C_{p})).$
By Brauer’s theorem on $l_{k}(G)$, we have $l_{k}(G)=\#C(G)+\frac{\#H-\#C(G)}{p}$ and $l_{k}(N_{G}(C_{p}))=l_{k}(C(G)\times C_{p})=\#C(G)$. Therefore, the number of isomorphic classes of indecomposable $kG$-modules coincides #$\#\mathrm{Conj}(G)$.
If $G$ is abelian, then #$\#\mathrm{Conj}(G)=\#G=p\#H$, $l_{k}(G)=\#H$ and $N_{G}(C_{p})=G$. Then easy computation shows the statement again. ∎
By Corollary 2.11 and Proposition 2.14, to prove the main theorem, it suffices to show that under the assumption of the main theorem, if there exists a crepant resolution, then
$S(F_{G})=\#\mathrm{Conj}(G).$
Note that even though we use a version of wild McKay correspondence over finite fields, we can consider the main theorem over appropriate finite fields instead, such that the $G$-actions, quotient singularties and resolutions are obtained by base change. In the remaining parts, we always assume that the order of the finite field $k=\mathbb{F}_{q}$ is large enough, such that it is a representation-theoretic splitting field of $G$ in characteristic $p$; in other words, not only Euler characteristic, but also the representation-theoretic invariants in the main theorem remain unchanged under base change to $\overline{k}$.
3 Abelian case of the main theorem
The proof starts from the case when $G$ is abelian. Now let $G\cong H\times C_{p}$ for an abelian non-modular subgroup $H$. We first give an observation on structure of such $G$.
Lemma 3.1.
Let $G\cong H\times C_{p}$ be a finite subgroup of $\mathrm{GL}(n,k)$. Then the representation space $V$ has a decomposition $V\cong\bigoplus_{i=1}^{s}V_{i}$, such that each $V_{i}$ induces an indecomposable $C_{p}$-representation of dimension $11\leqslant d_{i}\leqslant p-$ 1 and $\forall h\in H$, $h$ acts as a scalar multiplication on $V_{i}$.
Proof.
For $G\cong H\times C_{p}$, we can assume that $H$ consists of diagonal matrices by similarity transformation. Denote the matrix of a generator of $C_{p}$ under such transformation by $a=(a_{ij})_{i,j}$. $\forall h=\mathrm{diag}(\alpha_{1},\dots,\alpha_{n})\in H$, from $ha=ah$ we have $\alpha_{i}a_{ij}=\alpha_{j}a_{ij}$ for any $i,j$. If any $h$ in $H$ is of the form $\mathrm{diag}(c,\dots,c)$, then the lemma is trivial. So we can assume that there exists some $\alpha_{i}\neq\alpha_{j}$, and therefore $a_{ij}=a_{ji}=0$. By changing the order of basis if necessary, this implies that $a$ is a block diagonal matrix. In other words, $V$, as a representation space of $G$, can be decomposed into direct sum of subrepresentations as eigenspaces of different eigenvalues of $h$. Repeat this process on each subspace in the obtained decomposition, and then we obtain $V=\bigoplus_{t}W_{t}$, such that $\forall h\in H$ and $\forall W_{t}$, $W_{t}$ is a subspace of an eigenspace of $h$. Decompose each $W_{t}$ into direct sum of indecomposable representations of $C_{p}$. Then every components is still a $G$-representation, and the desired decomposition $V=\bigoplus_{i=1}^{s}V_{i}$ is obtained. ∎
Remark 3.2.
Consequently, we can choose basis of each $V_{i}$ appropriately, such that $H$ consists of diagonal matrices, and a generator of $C_{p}$ has the form $\mathrm{diag}(J_{d_{1}}(1),J_{d_{2}}(1),\dots,J_{d_{s}}(1))$, where $J_{i}(\alpha)$ denotes the Jordan block of size $i$ and eigenvalue $\alpha$.
Then we want to compute $f_{G}$ for $G\cong H\times C_{p}$ a small finite subgroup of $\mathrm{SL}(n,k)$.
For an element $L\in A(G)$, forgetting the $G$-action on it, $L$ is a Galois extension obtained by choosing a Galois extension $L_{1}/K$ with Galois group $\mathrm{Gal}(L_{1}/K)\cong H$ and an Artin-Schreier extension $L_{2}/K$, such that $L/K=L_{1}L_{2}/K$. For $L_{2}$, it depends on the choice of an element $a\in K$ such that $L_{2}=K(\alpha)$ for $\alpha$ satisfying $\wp(\alpha)=\alpha^{p}-\alpha=a$. We may choose $a$ such that the fixed generator of $C_{p}$ maps $\alpha$ to $\alpha+1$. According to [yasuda2014cyclic], Proposition 2.5, for this $G$-étale $K$-algebra, up to isomorphism, the element $a$ for the Artin-Schreier extension should be chosen from
$K/\wp(K)\setminus\{0\}\cong\mathbb{F}_{q}/\wp(\mathbb{F}_{q})\oplus\bigoplus_{j>0,(p,j)=1}\mathbb{F}_{q}t^{-j}\setminus\{0\}.$
In other words, every element $L\in A(G)$ is determined by an element $L_{1}\in A(H)$ and an element $L_{2}\in A(C_{p})$ determined by $a\in K/\wp(K)\setminus\{0\}$. If we fix $L_{1}$ and the degree $-j\in\{n\in\mathbb{Z}\mid n\leqslant 0,n\notin p\mathbb{Z}_{-}\}$, then there are
$\left\{\begin{array}[]{ll}p-1&j=0\\ pq^{j-1-\left\lfloor\frac{j-1}{p}\right\rfloor}(q-1)&j>0\end{array}\right.$
choices of $C_{p}$-étale extension $L_{2}$ generated by the element $\alpha$ with $v_{L_{2}}(\alpha)=-j$. $L_{2}$ is unramified if and only if $j=0$.
Fix $L\in A(G)$ and continue using notations $L_{1},L_{2},j$. We want to compute the value of $\mathbf{v}(L)$.
Lemma 3.3.
$\mathbf{v}(L)=\mathbf{v}(L/L_{1}^{\mathrm{ur}}),$ where $\mathbf{v}(L/L_{1}^{\mathrm{ur}})$ is the value of $\mathbf{v}$ funtion defined for group $\overline{G}:=G/\mathrm{Gal}(L_{1}^{\mathrm{ur}}/K)$ over $L_{1}^{\mathrm{ur}}\cong\mathbb{F}_{q^{f}}((t))$, and $f$ is the residue degree of $L_{1}/K$.
Proof.
Denote the dual basis of $T=\sum\mathcal{O}_{K}x_{i}$ in $\mathrm{Hom}_{\mathcal{O}_{K}}(T,\mathcal{O}_{L})$ by $x_{1}^{*},\dots,x_{n}^{*}$. Then by [yasuda2017toward], Proposition 6.3, $\mathrm{Hom}_{\mathcal{O}_{K}}^{G}(T,\mathcal{O}_{L})$ is a free $\mathcal{O}_{K}$-module of rank $n$. If $(\sum c_{ij}x_{j}^{*})_{1\leqslant i\leqslant n}$ is an $\mathcal{O}_{K}$-basis of $\mathrm{Hom}_{\mathcal{O}_{K}}^{G}(T,\mathcal{O}_{L})$, then $\mathbf{v}(L)=\frac{1}{\#G}\mathrm{length}_{\mathcal{O}_{K}}\frac{\mathcal{O}_{L}}{(\mathrm{det}(c_{ij}))}=\frac{v_{L}(\mathrm{det}(c_{ij}))f_{L/K}}{e_{L/K}f_{L/K}}=\frac{v_{L}(\mathrm{det}(c_{ij}))}{e_{L/K}}$.
By abuse of notation, let $(\sum a_{ij}x_{j}^{*})_{1\leqslant i\leqslant n}$ be an $\mathcal{O}_{L_{1}^{\mathrm{ur}}}$-basis of the module $\mathrm{Hom}_{\mathcal{O}_{L_{1}^{\mathrm{ur}}}}^{\overline{G}}(T,\mathcal{O}_{L})$. On the other hand, there exists an element $g\in\mathrm{Gal}(L_{1}^{\mathrm{ur}}/K)$ of order $f$ and $c\in\mathbb{F}_{q^{f}}\hookrightarrow L_{1}^{\mathrm{ur}}$ such that
$g(c)=\zeta_{f}c.$
The corresponding matrix of $g$ in $G$ has the form $\mathrm{diag}(\zeta_{f}^{a_{1}},\dots,\zeta_{f}^{a_{n}})$ ($0\leqslant a_{i}<f$ for each $i$). Then $(c^{a_{1}}x_{1}^{*},\dots,c^{a_{n}}x_{n}^{*})$ forms an $\mathcal{O}_{K}$-basis of the module $\mathrm{Hom}_{\mathcal{O}_{K}}^{\mathrm{Gal}(L_{1}^{\mathrm{ur}}/K)}(T,\mathcal{O}_{L_{1}^{\mathrm{ur}}})$. What is more, $(\sum a_{ij}c^{a_{j}}x_{j}^{*})_{1\leqslant i\leqslant n}$ is an $\mathcal{O}_{K}$-basis of $\mathrm{Hom}_{\mathcal{O}_{K}}^{\overline{G}}(T,\mathcal{O}_{L})$. Note that $v_{L}(c)=0$, and we have
$\mathbf{v}(L)=\frac{v_{L}(\mathrm{det}(c^{a_{j}}a_{ij}))}{e_{L/K}}=\frac{v_{L}(\mathrm{det}(a_{ij}))}{e_{L/L_{1}^{\mathrm{ur}}}}=\mathbf{v}(L/L_{1}^{\mathrm{ur}}).$
By Lemma 3.1 and Lemma 3.3, together with convertibility and additivity of $\mathbf{v}$ functions, we only need to consider the case when $L_{1}$ is a cyclic totally ramified extension and the representation is indecomposable.
Lemma 3.4.
Assume that $1<n\leqslant p$, and that $G$ is generated by $\sigma,\tau$ of the form
$\sigma=J_{n}(1),\tau=\mathrm{diag}(\zeta_{l},\dots,\zeta_{l}),$
and that $L_{1}$ is totally ramified with ramification index $l$. (We allow $G\subseteq\mathrm{GL}(n,k)$ here.)
Then when $l>1$,
$\mathbf{v}(L)=\frac{n}{l}+\sum_{i=1}^{n}\left\lceil\frac{(i-1)j}{p}-\frac{1}{l}\right\rceil.$
When $l=1$,
$\mathbf{v}(L)=\sum_{i=1}^{n}\left\lceil\frac{(i-1)j}{p}\right\rceil.$
Proof.
The case when $l=1$ has been computed in [yasuda2017toward], Example 6.8.
Now we assume $l>1$. Without loss of generality, we can choose the generator $\alpha$ of $L_{2}$ and a uniformizer $\beta$ of $\mathcal{O}_{L_{1}}$ such that
$\sigma(\alpha)=\alpha+1,\sigma(\beta)=\beta,\tau(\alpha)=\alpha,\tau(\beta)=\zeta_{l}\beta.$
We define formal binomial coefficients
(${\alpha\choose i}:=\frac{\alpha(\alpha-1)\dots(\alpha-i+1)}{i!}$
for $i=1,\dots,n-1$. Since $i<n\leqslant p$, they are well-defined. Additionally denote (${\alpha\choose 0}=1$. Then formal binomial coefficients satisfy
(${\alpha+1\choose i}={\alpha\choose i}+{\alpha\choose i-1}.$
Let $x_{1}^{*},\dots,x_{n}^{*}$ be the dual basis of $T=\sum\mathcal{O}_{K}x_{i}$ in $\mathrm{Hom}_{\mathcal{O}_{K}}(T,\mathcal{O}_{L})\hookrightarrow\mathrm{Hom}_{K}(\sum Kx_{i},L)$. Then the action of $\sigma$ and $\tau$ on $T$ is given by
$\sigma(x_{i})=\left\{\begin{array}[]{ll}x_{1}&i=1\\ x_{i-1}+x_{i}&i\neq 1\end{array}\right.,\tau(x_{i})=\zeta_{l}x_{i}.$
Then $\mathrm{Hom}^{G}_{K}(\sum Kx_{i},L)$ is generated by
$\begin{pmatrix}\beta x_{1}^{*}&\cdots&\beta x_{n}^{*}\end{pmatrix}\begin{pmatrix}0&0&\cdots&0&1\\ 0&0&\cdots&1&{\alpha\choose 1}\\ \vdots&\vdots&\iddots&\vdots&\vdots\\ 0&1&\cdots&\cdots&{\alpha\choose n-2}\\ 1&{\alpha\choose 1}&\cdots&{\alpha\choose n-2}&{\alpha\choose n-1}\\ \end{pmatrix}$
over $K$. Denote this basis by $(\phi_{1},\dots,\phi_{n})=\beta(x_{1}^{*},\dots,x_{n}^{*})\mathbf{A}$, and the matrix $\mathbf{A}$ has determinant ±$\pm 1$. If we take
$m_{i}:=\left\lceil\frac{-(i-1)v_{L}(\alpha)-v_{L}(\beta)}{e_{L/K}}\right\rceil=\left\lceil\frac{(i-1)j}{p}-\frac{1}{l}\right\rceil,$
then $\{t^{m_{1}}\phi_{1},\dots,t^{m_{n}}\phi_{n}\}$ forms an $\mathcal{O}_{K}$-basis of $\mathrm{Hom}_{\mathcal{O}_{K}}^{G}(T,\mathcal{O}_{L})$, and
$\mathbf{v}(L)=\frac{1}{e_{L/K}}v_{L}(\beta^{n}t^{m_{1}+\dots+m_{n}}\mathrm{det}\mathbf{A})$
shows the lemma.
Now we goes back to general case of the abelian group $G\cong H\times C_{p}$ and the $G$-étale $K$-algebra $L\in A(G)$.
Definition 3.5.
Consider the decomposition $V=\bigoplus_{i=1}^{s}V_{i}$ as in Lemma 3.1. Every indecomposable summand $V_{i}$ induces a Galois extension $N_{i}/L_{1}^{\mathrm{nr}}$, of which the tame ramification index is $l_{i}\geqslant 1$ and the Galois group is isomorphic to $C_{l_{i}}\times C_{p}$. We define the generalized upper shift number
$\mathrm{Sht}_{V}^{L}(j):=\sum_{i=1}^{s}\sum_{k=1}^{d_{i}}\left\lceil\frac{(k-1)j}{p}-\frac{1}{l_{i}}+\left\lfloor\frac{1}{l_{i}}\right\rfloor\right\rceil$
and the generalized age
$\mathrm{age}(L_{1}):=\sum_{i=1}^{s}d_{i}(\frac{1}{l_{i}}-\left\lfloor\frac{1}{l_{i}}\right\rfloor).$
Remark 3.6.
When $H=\{e\}$, $\mathrm{Sht}^{L}_{V}(j)=\mathrm{sht}_{V}(j)+n-s$, where $\mathrm{sht}_{V}$ is the shift number defined in [yasuda2014cyclic].
Corollary 3.7.
$\mathbf{v}(L)=\mathrm{age}(L_{1})+\mathrm{Sht}^{L}_{V}(j).$
Note that $l_{i}$ and $\mathrm{age}(L_{1})$ are totally determined by $L_{1}$. For the part of generalized upper shift numbers, it is determined by $L_{1}$ and $j$. We also have the following property by easy computation.
Proposition 3.8.
Denote $D_{V}:=\displaystyle\sum_{i=1}^{s}\frac{d_{i}(d_{i}-1)}{2}$. Then for each $j=ap+r$ ($a\geqslant 0,0<r<p$),
$\mathrm{Sht}^{L}_{V}(j)=aD_{V}+\mathrm{Sht}^{L}_{V}(r).$
With this property of $\mathrm{Sht}^{L}_{V}(j)$, we can compute $f_{G}$.
Proposition 3.9.
Let $G\cong H\times C_{p}$. If there exists a crepant resolution $Y\to X$, then $D_{V}=p$, and every $f_{G^{\prime}}$ with $G^{\prime}\cong H^{\prime}\times C_{p}\subseteq G$ converges as a polynomial, with $S(f_{G^{\prime}})=(p^{2}-1)S(f_{H^{\prime}}).$
Proof.
$f_{G}(q)$ $=\sum_{L\in A(G)}q^{n-\mathbf{v}(L)}$
$=\sum_{L_{1}\in A(H)}((p-1)q^{n-\mathrm{age}(L_{1})}$
$+\sum_{j>0,(p,j)=1}pq^{j-1-\left\lfloor\frac{j-1}{p}\right\rfloor}(q-1)q^{n-(\mathrm{age}(L_{1})+\mathrm{Sht}^{L}_{V}(j))})$
$=\sum_{L_{1}\in A(H)}q^{n-\mathrm{age}(L_{1})}(p-1$
$+p(1-q^{-1})\sum_{a=0}^{\infty}\sum_{r=1}^{p-1}q^{ap+r-a-(aD_{V}+\mathrm{Sht}^{L}_{V}(r))})$
$=\sum_{L_{1}\in A(H)}q^{n-\mathrm{age}(L_{1})}(p-1$
$+p(1-q^{-1})\sum_{r=1}^{p-1}q^{r-\mathrm{Sht}^{L}_{V}(r)}\sum_{a=0}^{\infty}q^{a(p-1-D_{V})}).$
Therefore, $f_{G}(q)$ converges if and only if $D_{V}\geqslant p$; if so, we can write
$f_{G}(q)$
$=$ $\sum_{L_{1}\in A(H)}q^{n-\mathrm{age}(L_{1})}(p-1+p\frac{1-q^{-1}}{1-q^{p-1-D_{V}}}\sum_{r=1}^{p-1}q^{r-\mathrm{Sht}^{L}_{V}(r)}).$
If this rational function is furthermore a polynomial, then $D_{V}=p$, and the value of $f_{G}$ at $q=1$ is $S(f_{G})=f_{G}(1)=\#A(H)(p-1+p(p-1))=(p^{2}-1)S(f_{H})$.
Since $D_{V}$ is equal for all $G^{\prime}\cong H^{\prime}\times C_{p}\subseteq G$, we complete the proof of this proposition.
Remark 3.10.
As a byproduct of this proposition, if $D_{V}\geqslant p$, then the quotient singularity $X$ is log terminal because of the convergence of stringy motif. The reader can refer to [yasuda2014cyclic] for details in $p$-cyclic McKay correspondence.
Corollary 3.11.
Main theorem holds when $G$ is abelian.
Proof.
By Proposition 3.9, if there exists a crepant resolution, then
$S(F_{G})$ $=\frac{\sum_{G^{\prime}\subseteq G}S(f_{G})}{\#G}=\frac{1}{\#G}\sum_{H^{\prime}\subseteq H}(S(f_{H^{\prime}})+S(f_{H^{\prime}\times C_{p}}))$
$=\frac{p^{2}(\#H)^{2}}{\#G}=\#G.$
4 Non-abelian case of the main theorem
To study non-abelian case, we need the following lemma.
Lemma 4.1.
Let $G$ be a non-abelian group with a semidirect product structure $H\rtimes C_{p}$. Then there does not exist any Galois extension $L/K$ such that $\mathrm{Gal}(L/K)\cong G$. In particular, $f_{G}=0$.
Proof.
If there exists such a Galois extension $L/K$, then $L^{C(G)}/K$ is a Galois extension with Galois group isomorphic to $G/C(G)\cong(H/C(G))\rtimes C_{p}$. Therefore, we can assume $C(G)=\{e\}$ without loss of generality.
Assume that $G$ is the Galois group of a Galois extension $L/K$. Consider the ramification groups
$G=G_{-1}\supseteq G_{0}\supseteq G_{1}\supseteq\dots.$
By classical results (one can refer to [serre2013local], IV, Corollary 4), $G_{0}$, as a normal subgroup of $G$, is a semidirect product of a normal $p$-subgroup and a non-modular cyclic group. If $G_{0}$ contains an element of order $p$, then $G_{0}$ has to be the whole $G$ since $G_{0}\triangleleft G$, which contradicts to the semidirect product structure of $G_{0}$. Hence $C_{p}$ is not contained in $G_{0}$. Since $G_{0}=\{g\in G\mid g\text{ acts trivially on }\mathrm{res}(L)\}$, $C_{p}$ acts effectively on $\mathrm{res}(L)$.
Hereby, for $L^{H}/K$, as an Artin-Schreier extension which is either unramified or totally ramified, its residue degree $f_{L^{H}/K}>1$. Consequently, $L^{H}/K$ is unramified. Therefore, if we write the local field $L\cong\mathrm{res}(L)((t_{L}))$, there exists an element $g\in G$ of order $p$ acting trivially on $t_{L}$. We assume that $C_{p}$ is generated by such $g$ without loss of generality.
Take an element $\sigma$ of prime order $l$ ($l\neq p$) from $H$. Then $L/L^{<\sigma>}$ is studied by Kummer theory. If $L/L^{<\sigma>}$ is unramified, then we can similarly assume that $\sigma$ acts trivially on $t_{L}$, and thus $\sigma$ has to commute with $C_{p}$, since Galois groups of extensions of finite fields are abelian. If $L/L^{<\sigma>}$ is totally ramified, then $\sigma$ acts trivially on $\mathrm{res}(L)$, and hence also commutes with $C_{p}$.
Either way, we can obtain a nontrivial element in $H$ that commutes with $C_{p}$ and thus lies in $C(G)$. Then the lemma is shown by contradiction, considering the assumption that $C(G)=\{e\}$. ∎
Remark 4.2.
Lemma 4.1 only holds in equal characteristic. In mixed characteristic, for example, the splitting field of $x^{13}+3$ over $\mathbb{Q}_{3}$ has non-abelian Galois group isomorphic to $C_{13}\rtimes C_{3}$ ([lmfdb:3.13.12.1]).
Proof of main theorem.
Since abelian case has been shown in Corollary 3.11, we only need to prove it for non-abelian group $G\cong H\rtimes C_{p}$ now. In this case, $C(G)\subseteq H$ and $N_{G}(C_{p})=C(G)\times C_{p}$.
For the polynomial $F_{G}$, we first write
$S(F_{G})=S(\sum_{[G^{\prime}]\subseteq H}\frac{f_{G^{\prime}}}{\#N_{G}(G^{\prime})}+\sum_{[G^{\prime}]:\text{modular}}\frac{f_{G^{\prime}}}{\#N_{G}(G^{\prime})}).$
Note that for $G^{\prime}\subseteq H$, if $N_{G}(G^{\prime})=G$, then the class $[G^{\prime}]$ contains $1$ element; if $N_{G}(G^{\prime})\neq G$, then $N_{G}(G^{\prime})=H$ and the class $[G^{\prime}]$ contains $p$ elements. Either way, if we write the sum over $G^{\prime}\subseteq H$, then the denominator is #$\#G=p\#H$. By Proposition 2.12,
$S(\sum_{[G^{\prime}]\subseteq H}\frac{f_{G^{\prime}}}{\#N_{G}(G^{\prime})})=S(\sum_{G^{\prime}\subseteq H}\frac{f_{G^{\prime}}}{\#G})=\frac{(\#H)^{2}}{\#G}=\frac{\#H}{p}.$
By Lemma 4.1, for modular subgroups $G^{\prime}\subseteq G$, $f_{G^{\prime}}=0$ unless $G^{\prime}\subseteq C(G)\times C_{p}$. Therefore, by Proposition 3.9, assuming existence of crepant resolutions,
$S(\sum_{[G^{\prime}]:\text{modular}}\frac{f_{G^{\prime}}}{\#N_{G}(G^{\prime})})=S(\sum_{C_{p}\subseteq G^{\prime}\subseteq C(G)\times C_{p}}\frac{f_{G^{\prime}}}{p\#C(G)})$
$=\sum_{H^{\prime}\subseteq C(G)}\frac{(p^{2}-1)S(f_{H^{\prime}})}{p\#C(G)}=(p-\frac{1}{p})\#C(G).$
Hence $S(F_{G})=p\#C(G)+\frac{\#H-\#C(G)}{p}=\#\mathrm{Conj}(G)$. ∎
5 Computation of Euler characteristic
The idea of proof of main theorem can also be applied for computation of Euler characteristic in other cases: in [yamamoto2022mass], Yamamoto computed a series of Euler characteristic, when the group has a semidirect product structure of a non-modular abelian normal subgroup and the symmetric group $S_{3}$, in characteristic $3$ and dimension $3$. In this section, we introduce examples where the $p$-Sylow subgroup of $G$ is not cyclic (equivalently, the group $G$ is of infinite representation type in characteristic $p$; it is a result by Higman[higman1954indecomposable]), and see that Euler characteristic can still be computed via wild McKay correspondence. To compute $\mathbf{v}$ function in the examples, we need the theorem below.
Theorem 5.1 ([wood2015mass], Theorem 4.8).
If $G$ acts on $V$ by permutation, then $\mathbf{v}(L)=\frac{1}{2}\mathbf{a}(L)$, where
$\mathbf{a}(L):=\sum_{i=0}^{\infty}\frac{\mathrm{codim}(k^{n})^{G_{i}}}{(G_{0}:G_{i})}$
is the Artin conductor of this permutation representation.
In the remaining part of this section, $k$ is a finite field of order $q>2$ and characteristic $2$, $A_{4}$ is the alternating group with permutation action on $\mathbb{A}_{k}^{4}$, and $C_{2}^{2}=\{e,(12)(34),(13)(24),(14)(23)\}$ is the normal subgroup consisting of modular elements in $A_{4}$, also acting on $\mathbb{A}^{4}$ by permutations.
Proposition 5.2.
For an element $L\in A(C_{2}^{2})$, the corresponding extension is determined by composition of two Artin-Schreier extensions, which are given by two different elements
$a,b\in k/\wp(k)\oplus\bigoplus_{j>0,(2,j)=1}kt^{-j}\setminus\{0\}.$
Denote $v_{K}(a)=-j$ and $v_{K}(b)=-k$, and assume that $j\leqslant k$. Then
$\mathbf{v}(L)=\left\{\begin{array}[]{ll}\displaystyle\frac{1}{2}(j+1)+k+1&j>0\\ \displaystyle k+1&j=0\end{array}\right.$
Proof.
By [wu2010ramification], Theorem 3.11, if $L$ is totally ramified, then the ramification groups of $C_{2}^{2}$ are
$G_{0}=\dots=G_{j}=C_{2}^{2},G_{j+1}=\dots=G_{j+2(k-j)}=C_{2},G_{j+2(k-j)+1}=\{e\}.$
Therefore, by Theorem 5.1, when $j>0$, $\mathbf{v}(L)=\frac{3}{2}(j+1)+\frac{1}{2}(2k-2j)=\frac{1}{2}(j+1)+k+1$; when $j=0$, the ramification groups are determined only by the Artin-Schreier extension corresponding to $b$, and $\mathbf{v}(L)=k+1$. ∎
Proposition 5.3.
For an element $L\in A(A_{4})$, the corresponding extension is determined by $L_{1}=K(\gamma)\in A(C_{3})$ and an Artin-Schreier extension over $L_{1}$ given by an element
$a\in\left\{\begin{array}[]{ll}\displaystyle\bigoplus_{j>0,(2,j)=1}(\gamma k\oplus\gamma^{2}k)t^{-j}&L_{1}/K\text{ is unramified}\\ \displaystyle\bigoplus_{j>0,(6,j)=1}k\gamma^{-j}&L_{1}/K\text{ is ramified}\end{array}\right.$
Denote $v_{L_{1}}(a)=-j$, and then
$\mathbf{v}(L)=\left\{\begin{array}[]{ll}\displaystyle\frac{3}{2}(j+1)&L_{1}/K\text{ is unramified}\\ \displaystyle\frac{1}{2}(j+3)&L_{1}/K\text{ is ramified}\end{array}\right.$
Proof.
Assume that $L=L_{1}(\alpha,\beta)$, where $L_{1}(\alpha)$ and $L_{1}(\beta)$ are Artin-Schreier extensions corresponding to $a,b\in L_{1}$. Then all the nontrivial intermediate fields of $L/L_{1}$ are $L_{1}(\alpha),L_{1}(\beta)$ and $L_{1}(\alpha+\beta)$. Then any element of order $3$ in $\mathrm{Gal}(L/K)\cong A_{4}$ should give a cyclic permutation on $\alpha,\beta$ and $\alpha+\beta$.
In other words, $L/L_{1}$ is determined by an element $a\in L_{1}$ such that for a generator $\tau\in\mathrm{Gal}(L_{1}/K)$, $a,\tau(a)$ and $\tau^{2}(a)$ are distinct and $\tau^{2}(a)=\tau(a)+a$. Assume that $L_{1}=K(\gamma)$ and $\tau(\gamma)=\zeta_{3}\gamma$, where $\gamma^{3}$ is in $\mathrm{res}(K)$ if $L_{1}/K$ is unramified, while $\gamma$ a uniformizer of $\mathcal{O}_{L_{1}}$ if $L_{1}/K$ is totally ramified. Then computation shows that $a$ can be taken as stated.
Then we compute value of $\mathbf{v}$ function. If $L_{1}/K$ is unramified, then the ramification groups are determined by $C_{2}^{2}$ part as in proof of Proposition 5.2, where $v_{L_{1}}(a)=v_{L_{1}}(\tau(a))=v_{L_{1}}(\tau^{2}(a))=-j$, hence $\mathbf{v}(L)=\frac{3}{2}(j+1)$; if $L_{1}/K$ is ramified, then the ramification groups are
$G_{0}=A_{4},G_{1}=\dots=G_{j}=C_{2}^{2},G_{j+1}=\{e\},$
and then by Theorem 5.1, $\mathbf{v}(L)=\frac{3}{2}+\frac{1}{2}j=\frac{1}{2}(j+3)$. ∎
Proposition 5.4.
$f_{C_{2}^{2}}=10q^{3}+4q^{2},f_{A_{4}}=32q^{3}+32q^{2}.$
Proof.
For $L\in A(C_{2}^{2})$, it is determined by $a,b$ as stated in Proposition 5.2. However, considering all the three nontrivial intermediate fields of $L/K$, we know that any two elements from $\{a,b,a+b\}$ give the same extension. We may furthermore assume that $v_{L}(a)\geqslant v_{L}(b)=v_{L}(a+b)$. On the other hand, for each extension, there are $3!=6$ ways to equip it with a $C_{2}^{2}$-action by mapping nontrivial elements in $C_{2}^{2}$ to nontrivial elements of $\mathrm{Gal}(L/K)$.
Therefore, there are
$\frac{6\times 2q^{\frac{k-1}{2}}(q-1)}{2}$
choices for $L\in A(C_{2}^{2})$ with $j=0,k>0$,
$\frac{6\times(2q^{\frac{j-1}{2}}(q-1))(2q^{\frac{j-1}{2}}(q-2))}{6}$
choices for $L\in A(C_{2}^{2})$ with $j=k>0$, and
$\frac{6\times(2q^{\frac{j-1}{2}}(q-1))(2q^{\frac{k-1}{2}}(q-1))}{2}$
choices for $L\in A(C_{2}^{2})$ with $0<j<k$.
Consequently,
$f_{C_{2}^{2}}$ $=\sum_{k>0,(2,k)=1}6q^{\frac{k-1}{2}}(q-1)q^{4-(k+1)}$
$+\sum_{j>0,(2,j)=1}(2q^{\frac{j-1}{2}}(q-1))(2q^{\frac{j-1}{2}}(q-2))q^{4-\frac{3}{2}(j+1)}$
$+\sum_{j>0,(2,j)=1}\sum_{k>j,(2,k)=1}3(2q^{\frac{j-1}{2}}(q-1))(2q^{\frac{k-1}{2}}(q-1))q^{4-\frac{1}{2}(j+1)-(k+1)}$
$=6q^{4}(q-1)\sum_{i=0}^{\infty}q^{i}q^{-2i-2}$
$+4(q-1)(q-2)q^{4}\sum_{i=0}^{\infty}q^{i}q^{i}q^{-3i-3}$
$+12(q-1)^{2}q^{4}\sum_{r=0}^{\infty}\sum_{s=1}^{\infty}q^{r}q^{r+s}q^{-r-1-2r-2s-2}$
$=6q^{2}(q-1)\frac{1}{1-q^{-1}}+4(q-1)(q-2)q\frac{1}{1-q^{-1}}+12(q-1)^{2}q\frac{1}{1-q^{-1}}\frac{q^{-1}}{1-q^{-1}}$
$=6q^{3}+4q^{2}(q-2)+12q^{2}=10q^{3}+4q^{2}.$
We use the same way to compute $f_{A_{4}}$. For $L\in A(A_{4})$, it is determined by $L_{1}\in A(C_{3})$ and $a\in L_{1}$ as in Proposition 5.3. Since $L_{1}/K$ is a Kummer extension, we have $1$ choice for unramified one and $3$ choices for totally ramified ones (ignoring $C_{3}$-action on it). By abuse of notation, let $\tau$ be an element in $\mathrm{Gal}(L/K)$ of order $3$, whose restrction on $L_{1}$ generates $\mathrm{Gal}(L_{1}/K)$. For $L/L_{1}$, on one hand, $a,\tau(a),\tau^{2}(a)$ give the same extension; on the other hand, there are $3!=6$ ways to equip the extension with a $C_{2}^{2}$-action. Finally, whenever the $C_{2}^{2}$-extension is determined, one can choose an element of order $3$ in $A_{4}$ from a specific conjugacy class such that it acts on $L$ as $\tau$, which contains $4$ choices.
Therefore, there are
$\frac{6\times 4\times(q^{2})^{\frac{j-1}{2}}(q^{2}-1)}{3}$
choices for $L\in A(A_{4})$ with $L_{1}$ unramified, $v_{L_{1}}(a)=-j$, and
$3\times\frac{6\times 4\times q^{j-1-\left\lfloor\frac{j-1}{2}\right\rfloor-\left\lfloor\frac{j-1}{3}\right\rfloor+\left\lfloor\frac{j-1}{6}\right\rfloor}(q-1)}{3}$
choices for $L\in A(A_{4})$ with $L_{1}$ totally ramified, $v_{L_{1}}(a)=-j$.
Consequently,
$f_{A_{4}}$ $=\sum_{j>0,(2,j)=1}8q^{j-1}(q^{2}-1)q^{4-\frac{3}{2}(j+1)}$
$+\sum_{j>0,(6,j)=1}24q^{j-1-\left\lfloor\frac{j-1}{2}\right\rfloor-\left\lfloor\frac{j-1}{3}\right\rfloor+\left\lfloor\frac{j-1}{6}\right\rfloor}(q-1)q^{4-\frac{1}{2}(j+3)}$
$=8(q^{2}-1)q^{3}\sum_{i=0}^{\infty}q^{2i+1}q^{-3i-3}$
$+24(q-1)q^{3}\sum_{r=0}^{\infty}(q^{2r+1}q^{-3r-2}+q^{2r+2}q^{-3r-4})$
$=8(q^{2}-1)q\frac{1}{1-q^{-1}}+24(q-1)q(\frac{q}{1-q^{-1}}+\frac{1}{1-q^{-1}})$
$=8(q+1)q^{2}+24(q^{3}+q^{2})=32q^{3}+32q^{2}.$
Corollary 5.5.
If the quotient singularity $X_{1}:=\mathbb{A}^{4}/C_{2}^{2}$ ($resp.$ $X_{2}:=\mathbb{A}^{4}/A_{4}$) has a crepant resolution $Y_{1}\to X_{1}$ ($resp.$ $Y_{2}\to X_{2}$), then $e(Y_{1})=6$ ($resp.$ $e(Y_{2})=10$).
Proof.
$S(F_{C_{2}^{2}})$ $=\frac{1}{4}(S(f_{\{e\}})+3S(f_{C_{2}})+S(f_{C_{2}^{2}}))=\frac{1}{4}(1+9+14)=6,$
$S(F_{A_{4}})$ $=\frac{S(f_{\{e\})})}{12}+\frac{S(f_{C_{2}})}{4}+\frac{S(f_{C_{3}})}{3}+\frac{S(f_{C_{2}^{2}})}{12}+\frac{S(f_{A_{4}})}{12}$
$=\frac{1}{12}+\frac{3}{4}+\frac{8}{3}+\frac{14}{12}+\frac{64}{12}=10.$
Remark 5.6.
Although the author does not know if $X_{1}$ has a crepant resolution, $X_{2}$ does have a crepant resolution with Euler characteristic $10$, as constructed in [fan2023crepant].
Remark 5.7.
We can furthermore compute that $F_{C_{2}^{2}}=q^{4}+4q^{3}+q^{2}$ and $F_{A_{4}}=q^{4}+6q^{3}+3q^{2}$, which can be seen as a realization of stringy motives for quotient singularities $X_{1}$ and $X_{2}$ over $\mathbb{F}_{q}$, from the perspective of motivic wild McKay correspondence. This coincides the construction in [fan2023crepant], where the class of crepant resolution of $\mathbb{A}^{4}/A_{4}$ in the (modified) Grothendieck ring is $\mathbb{L}^{4}+6\mathbb{L}^{3}+3\mathbb{L}^{2}$.